Suppose an indecisive man starts out from home and walks 1 mi east,then 1/2 mi west, then 1/4 mi east, then 1/8 mi west, and so on. Relative to his home, approximately where would he end up?

I have no clue..please help.
Thanks

Answer by stanbon(75887)

Suppose an indecisive man starts out from home and walks 1 mi east,then 1/2 mi west, then 1/4 mi east, then 1/8 mi west, and so on. Relative to his home, approximately where would he end up?
-----------------
Total lateral distance: 1-(1/2)+(1/4)-(1/8)+(1/16)-(1/32)....
= [1 + 1/4 + 1/16 + 1/64... ] - [1/2 + 1/8 + 1/32]
---
Using the sum of a geometric series you get:
= (1)[1/(1-(1/4)] - (1/2)[1-(1/4)]
---
= 1/(3/4) - (1/2)/(3/4)
---
= 4/3 - 2/3
---
= 2/3
-----------------------------
Cheers,
Stan H.
========================